# thanks: Contributed equallythanks: Contributed equally

### v1

@@ Abstract
Continuous-time quantum walks (CTQWs) on dynamic graphs, referred to as dynamic CTQWs, are a recently introduced universal model of computation that offers a new paradigm in which to envision quantum algorithms. In this work we develop a mapping from dynamic CTQWs to the gate model of computation in the form of an algorithm to convert arbitrary single edge walks and single self loop walks, which are the fundamental building blocks of dynamic CTQWs, to their circuit model counterparts. We use this mapping to introduce an arbitrary quantum state preparation framework based on dynamic CTQWs. Our approach utilizes global information about the target state, relates state preparation to finding the optimal path in a graph, and leads to optimizations in the reduction of controls that are not as obvious in other approaches. Interestingly, classical optimization problems such as the minimal hitting set, minimum spanning tree, and shortest Hamiltonian path problems arise in our framework. We test our methods against uniformly controlled rotations methods, used by Qiskit, and find ours requires fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
@@ I Introduction
However, the opposite conversion from CTQWs to the gate model is less studied. Since CTQWs on dynamic graphs have yet to be implemented on hardware, converting them to the gate model is necessary to simulate them on existing hardware. Additionally, developing an algorithm for such a conversion offers an alternative way of thinking about circuit design, which may result in simpler circuits, compiler optimization techniques, and new ansatzes for QAOA [21, 22, 23, 24], MA-QAOA [25, 26, 27] or VQE [28, 29, 30], where the trainable parameters correspond to dynamic graph propagation times.
To this end, we develop an algorithm that converts CTQWs on single edge graphs and single self-loop graphs, which can serve as a basis for arbitrary graphs, to a sequence of gates in the circuit model. This conversion algorithm is used as a foundation for a new deterministic arbitrary quantum state preparation (QSP) method, where the CX count of the resulting circuit is linear in the number of $m$ non-zero amplitude computational basis states that comprise the target state and the number of qubits $n$. Fig. 1 gives a general picture of the quantum walks QSP framework introduced in this paper. Our framework does not require ancillas.
Figure 1: This figure lays out the framework for QSP via quantum walks. Alternating single edge quantum walks with self-loop quantum walks, allows arbitrary QSP. We provide conversion methods to construct the gate based circuit. Control reduction is a method for reducing controls that we introduce in this paper.
Our method builds on the idea of exploiting sparsity to increase the efficiency of QSP techniques by approaching the problem from the perspective of dynamic quantum walks. As such, our method is best suited for the asymptotically sparse states ($m=\mathcal{O}(poly(n))$), but can, in principle, also work for the asymptotically dense states ($m=\mathcal{O}(2^{n})$). The walk framework that we present here is flexible, scalable, and intuitive, building upon well-established graph-based algorithms to efficiently prepare arbitrary quantum states.
This paper is organized as follows. First, we introduce necessary background information and previous works in QSP in Sec. II. We then introduce an algorithm that converts CTQWs on single edge dynamic graphs and single self loop graphs to the circuit model in Sec. III. In Sec. IV, we introduce a deterministic state preparation algorithm and analyze its circuit complexity. We compare our state preparation methods to the uniformly controlled rotation method [5, 31] used by Qiskit in Sec. V. Finally, we discuss future work in Sec. VI.
@@ II.2 Prior State Preparation Results
Due to its fundamental role in quantum algorithms [32, 33, 34], quantum state preparation is an active area of research within the space of quantum computation. The resource overhead of general QSP problems for arbitrary states is known to be $\mathcal{O}(2^{n})$ in the number of CX gates [5, 31]. Within this limit, there are different techniques utilizing different resources, which can be useful depending on the context. For instance, if depth is more of a concern than space, Ref. [35] provides an $\mathcal{O}(n^{2})$ depth divide-and-conquer method to prepare arbitrary states at the cost of needing $\mathcal{O}(2^{n})$ ancillary qubits. For arbitrary QSP, the state of the art deterministic protocol achieves $\mathcal{O}(\frac{23}{24}2^{n})$ CX scaling on even numbers of qubits [36]. The downside of this method is that it relies on the Schmidt decomposition of the $n$-qubit target state, a computationally expensive task, which has a classical run time of $\mathcal{O}(2^{3n/2})$ [37].
@@
The state of the art method for preparing arbitrary sparse states is detailed in [6], which produces circuits with $\mathcal{O}(nm)$ CX gates and runs in $\mathcal{O}(nm^{2}\log(m))$ time classically. This type of method is attractive for real-world applications of quantum computers because they can take advantage of sparsity while still remaining agnostic to particular details of the state. Another method using decision trees [43] has $\mathcal{O}(nm)$ CX gates, but it requires an ancilla qubit. The method in Ref. [44] also achieves $\mathcal{O}(nm)$ CX gates. However, this method is based on Householder reflections and is more complicated than the one introduced in Ref. [6].
@@ III Conversion from CTQW on dynamic graphs to the Circuit Model
A standard basis for any adjacency matrix is given by the set of adjacency matrices corresponding to all single edge walks and all single self-loop walks. We introduce the conversion methods that can construct the gate-based representations for these walks on $n$ qubits. Throughout this paper, the considered states are represented in the standard computational basis set.
III.1 Single Edge CTQW to Circuit Model
The adjacency matrix for a single edge graph connecting basis states |$\ket{j}$ and |$\ket{k}$ is given by
@@ III.1 Single Edge CTQW to Circuit Model
If the Hamming distance between |$\ket{j}$ and |$\ket{k}$ is greater than $1$, using the CRx($2t$) gate is still possible, but requires some extra overhead. First apply a sequence of CX gates where the control is any bit in which |$\ket{j}$ and |$\ket{k}$ differ and the targets are the other bits in which |$\ket{j}$ and |$\ket{k}$ differ. Then apply the CRx($2t$) gate as before and apply the reverse of the conjugating CX sequence. An example conversion of a single edge walk to gates is shown in Fig. 3. The first CX gate transforms the two basis states such that their Hamming distance becomes equal to 1. The CRx creates the superposition and the final CX restores the original basis states.
@@
Figure 3: When a single edge walk is between states with non-unit Hamming distance from each other, CX gates might be required in addition to the CRx gate.
@@ III.3 Universal Computation
As it was mentioned in the Introduction, CTQWs on dynamic graphs are known to be universal [17]. However, the authors of Ref. [17] use walks on arbitrary graphs in their proof of universality. In this section we prove that using only single edge and single self-loop walks is sufficient to decompose an arbitrary unitary.
@@
An arbitrary $d\times d$ unitary can be decomposed into a series of single self-loop and single edge CTQWs.
@@
As it was shown in Sections III.1 Eq. (3) and III.2 Eq. (6), a single edge CTQW corresponds to Rx’, and a sequence of two single self-loop CTQWs corresponds to W’ and Rz’. Thus, an arbitrary unitary can be decomposed into a series of single self-loop CTQWs and single edge CTQWs.
@@ IV Deterministic state preparation via quantum walks
The above conversions allow one to write single edge walk dynamics and single self-loops in terms of the circuit model. In this section, we describe how a series of such quantum walks can be used to prepare an arbitrary quantum state.
@@ IV.1 General Approach
Given the task of preparing a quantum state with $m<2^{n}$ non-zero amplitudes on $n$ qubits, we start by presenting a high-level description of the algorithm to accomplish this task:
@@
    Create a tree graph connecting the non-zero amplitudes, as shown in Fig. 4.
@@
    Starting from an arbitrarily chosen root, and following a graph traversal order, perform a self-loop walk for each encountered node and a single edge quantum walk for each encountered edge.
@@
    Convert the sequence of the quantum walks to the circuit representation.
Initially, the system starts with all amplitude in the root state, which can be constructed from the ground |$\ket{0}$ state simply by applying the X gates to the necessary qubits. The graph traversal order can be be determined by any graph traversal algorithm, such as Depth First Search or Breadth First Search.
An arbitrary amount of amplitude (absolute value) from the source can be transferred to the destination for any pair of the connected states corresponding to each single edge walk, and all non-zero amplitude states are connected transitively via a tree. Therefore, an arbitrary distribution of absolute values of amplitude (but not phases) can be established on the set of non-zero amplitude basis states via single edge walks.
Each self-loop walk on a basis state |$\ket{z_{i}}$ establishes the correct phase of the corresponding coefficient $c_{i}$, but does not change the absolute value of it. Therefore, performing self-loop walks on each non-zero amplitude basis state establishes the correct phases of the corresponding coefficients, without interfering with the absolute values established by the single edge walk. Therefore, an arbitrary quantum state can be prepared with a sequence of single edge walks and self-loop walks, as described in the algorithm above.
In the context of the state preparation, CRz gates can easily be used instead of CP gates to implement the self-loop walks, since every single edge walk transfers the amplitude to a new basis state with zero amplitude. Thus, as described in the previous section, the same basis state can be used to establish the correct phase on the source state for the single edge walk without additional CX conjugations. However, this will not apply to the leaf nodes (degree one vertices) in the tree, which will require another zero-amplitude state to interact with.
@@ IV.2 Control Reduction
Let $n$ denote the number of qubits and $S=\{\ket{z_{1}},\ket{z_{2}},...,\ket{z_{k}}\}$ denote the set of visited nodes. Suppose we wish to perform a single edge walk from |$\ket{z_{j}}\in S$ to |$\ket{z_{\ell}}\notin S$ and |$\ket{z_{j}}$ and |$\ket{z_{\ell}}$ have a Hamming distance of 1. Let $b$ denote the qubit where |$\ket{z_{j}}$ is different from |$\ket{z_{\ell}}$, i.e. $z_{j}[b]\neq z_{\ell}[b].$ Let $D=\{\{k\,|\,z_{i}[k]\neq z_{j}[k],k\neq b\}\,|\,z_{i}\in S\setminus z_{j}\}$ denote the set of differing bits of the visited nodes with $z_{j}$ excluding the qubit $b$. For the gate-based representation of a single edge walk from basis state |$\ket{z_{j}}$ to |$\ket{z_{\ell}}$, it is sufficient to control the CRz or CRx gate on any hitting set of $D$, where the values of controls are equal to the corresponding bits of $z_{j}$. A hitting set of $D$ is a collection of elements $h$ such that $h\cap d_{i}\neq\emptyset$ for all $d_{i}\in D$.
@@
When $z_{j}$ and $z_{\ell}$ have a Hamming distance greater than 1, we first have to update $S\rightarrow\tilde{S}$ with the conjugating CX gates, as described in Sec. III.1. Then we apply the previously described control reduction method to $\tilde{S}$.
In order to minimize the number of controls, one needs to minimize the size of the hitting set, i.e. solve the minimum hitting set problem, which is known to be NP-complete. As such, no polynomial algorithm is known to solve it exactly, but a number of heuristics exist that can provide good suboptimal solutions quickly [48, 49]. In practice, solving this problem for each walk drastically reduces the number of control qubits on many gates in the sequence (see Fig. 5).
@@
IV.3 Walk Order
Any tree constructed in the first step is, in principle, sufficient for state preparation but not all trees result in the same CX gate count. As mentioned in the previous section, the cost of a walk between a given pair of basis states depends on the Hamming distance between them. Therefore, connecting the basis states that are close to one another (in terms of their Hamming distance) minimizes the number of CX gates in the resulting circuit. Let us consider some of the more advantageous options.

### v2

# Efficient Sparse State Preparation via Quantum Walks

@@ Abstract
Continuous-time quantum walks (CTQWs) on dynamic graphs, referred to as dynamic CTQWs, are a recently introduced universal model of computation that offers a new paradigm in which to envision quantum algorithms. In this work we develop a mapping from dynamic CTQWs to the gate model of computation in the form of an algorithm to convert arbitrary single-edge walks and single self-loop walks, which are the fundamental building blocks of dynamic CTQWs, to their circuit model counterparts. We use this mapping to introduce an efficient sparse quantum state preparation framework based on dynamic CTQWs. Our approach utilizes global information about the target state, relates state preparation to finding the optimal path in a graph, and leads to optimizations in the reduction of controls that are not as obvious in other approaches. Interestingly, classical optimization problems such as the minimal hitting set, minimum spanning tree, and shortest Hamiltonian path problems arise in our framework. We test our methods against the state-of-the-art ancilla free sparse state preparation method and a uniformly controlled rotation method, which is used by Qiskit, and find ours requires fewer CX gates when the target state has a polynomial number of non-zero amplitudes.
@@ I Introduction
While prior work has focused on representing quantum gate operations as CTQWs, the opposite conversion – generating a quantum circuit from a CTQW – is less well-studied. Since CTQWs on dynamic graphs have yet to be natively implemented on hardware, and given recent progress in the development of gate-based quantum hardware [21, 22, 23, 24], converting CTQWs to the gate model is necessary to execute them on existing quantum hardware. Additionally, developing an algorithm for such a conversion offers an alternative way of thinking about circuit design, which may result in simpler circuits, compiler optimization techniques, and new ansatzes for QAOA [25, 26, 27, 28], MA-QAOA [29, 30, 31] or VQE [32, 33, 34], where the trainable parameters correspond to dynamic graph propagation times.
To this end, we develop an algorithm that converts CTQWs on single-edge graphs and single self-loop graphs, which can serve as a basis for arbitrary graphs, to a sequence of gates in the circuit model. This conversion algorithm is used as a foundation for a new deterministic arbitrary quantum state preparation (QSP) method, where the CX count of the resulting circuit is linear in the number of $m$ non-zero amplitude computational basis states that comprise the target state and the number of qubits $n$. Fig. 1 gives a general picture of the quantum walks QSP framework introduced in this paper. Our framework does not require ancillas.
Figure 1: Alternating single-edge and self-loop quantum walks yields a framework for arbitrary QSP. We provide conversion methods to construct the gate-based circuit. Control reduction is a method for reducing controls that we introduce in this paper.
Our method builds on the idea of exploiting sparsity to increase the efficiency of QSP techniques by approaching the problem from the perspective of dynamic quantum walks. As such, our method is best suited for the asymptotically sparse states ($m=\mathcal{O}(poly(n))$), but can, in principle, also work for the asymptotically dense states ($m=\mathcal{O}(2^{n})$). The walk framework that we present here is flexible, scalable, and intuitive, building upon well-established graph-based algorithms to efficiently prepare sparse quantum states.
This paper is organized as follows. First, we introduce necessary background information and previous works in QSP in Sec. II. We then introduce an algorithm that converts CTQWs on single-edge dynamic graphs and single self-loop graphs to the circuit model in Sec. III. In Sec. IV, we introduce a deterministic state preparation algorithm and analyze its circuit complexity. We compare our state preparation methods to the sparse state preparation method from [6] and to the uniformly controlled rotation method [5, 35] used by Qiskit in Sec. V. In both cases, we found that our algorithm requires fewer CX gates. Finally, we discuss future work in Sec. VI.
@@ II.2 Prior State Preparation Results
Due to its fundamental role in quantum algorithms [36, 37, 38], quantum state preparation is an active area of research within the space of quantum computation. The resource overhead of general QSP problems for arbitrary states is known to be $\mathcal{O}(2^{n})$ in the number of CX gates [5, 35]. Within this limit, there are different techniques utilizing different resources, which can be useful depending on the context. For instance, if depth is more of a concern than space, Ref. [39] provides an $\mathcal{O}(n^{2})$ depth divide-and-conquer method to prepare arbitrary states at the cost of needing $\mathcal{O}(2^{n})$ ancillary qubits. In Ref. [40], the ancilla overhead is improved to $\mathcal{O}(n)$. For arbitrary QSP, the state of the art deterministic protocol achieves $\mathcal{O}(\frac{23}{24}2^{n})$ CX scaling on even numbers of qubits [41]. The downside of this method is that it relies on the Schmidt decomposition of the $n$-qubit target state, a computationally expensive task, which has a classical run time of $\mathcal{O}(2^{3n/2})$ [42].
@@
The state of the art method for preparing arbitrary sparse states is detailed in [6], which produces circuits with $\mathcal{O}(nm)$ CX gates and runs in $\mathcal{O}(nm^{2}\log(m))$ time classically. This type of method is attractive for real-world applications of quantum computers because they can take advantage of sparsity while still remaining agnostic to particular details of the state. Other methods achieve this, but require ancillas [48, 49]. The method using decision trees [49] requires 1 ancilla qubit. The method in Ref. [50] also achieves $\mathcal{O}(nm)$ CX gates. However, this method is based on Householder reflections and is more complicated than the one introduced in Ref. [6].
@@ III Conversion from CTQW on dynamic graphs to the Circuit Model
A standard basis for any adjacency matrix is given by the set of adjacency matrices corresponding to all single-edge walks and all single self-loop walks. We introduce the conversion methods that can construct the gate-based representations for these walks on $n$ qubits. Throughout this paper, the considered states are represented in the standard computational basis set.
III.1 Single-Edge CTQW to Circuit Model
The adjacency matrix for a single-edge graph connecting basis states |$\ket{j}$ and |$\ket{k}$ is given by
@@ III.1 Single-Edge CTQW to Circuit Model
If the Hamming distance between |$\ket{j}$ and |$\ket{k}$ is greater than $1$, using the CRx($2t$) gate is still possible, but requires some extra overhead. First apply a sequence of CX gates where the control is any bit in which |$\ket{j}$ and |$\ket{k}$ differ and the targets are the other bits in which |$\ket{j}$ and |$\ket{k}$ differ. Then apply the CRx($2t$) gate as before and apply the reverse of the conjugating CX sequence. An example conversion of a single-edge walk to gates is shown in Fig. 3. The first CX gate transforms the two basis states such that their Hamming distance becomes equal to 1. The CRx creates the superposition and the final CX restores the original basis states.
@@
Figure 3: When a single-edge walk is between states with non-unit Hamming distance from each other, CX gates might be required in addition to the CRx gate.
@@ III.3 Universal Computation
As it was mentioned in the Introduction, CTQWs on dynamic graphs are known to be universal [17]. However, the authors of Ref. [17] use walks on arbitrary graphs in their proof of universality. In this section we prove that using only single-edge and single self-loop walks is sufficient to decompose an arbitrary unitary.
@@
An arbitrary $d\times d$ unitary can be decomposed into a series of single self-loop and single-edge CTQWs.
@@
As it was shown in Sections III.1 Eq. (3) and III.2 Eq. (6), a single-edge CTQW corresponds to Rx’, and a sequence of two single self-loop CTQWs corresponds to W’ and Rz’. Thus, an arbitrary unitary can be decomposed into a series of single self-loop CTQWs and single-edge CTQWs.
@@ IV Deterministic state preparation via quantum walks
The above conversions allow one to write single-edge walk dynamics and single self-loops in terms of the circuit model. In this section, we describe how a series of such quantum walks can be used to prepare an arbitrary quantum state.
@@ IV.1 General Approach
Given the task of preparing a quantum state with $m<2^{n}$ non-zero amplitudes on $n$ qubits, we start by presenting a high-level overview of the algorithm to accomplish this task:
@@
    Construct a tree graph connecting the non-zero amplitudes of the target state in some way. See Fig. 4 for an example.
@@
    Starting from some node, and following frontier-based graph traversal order (see Sec. IV.5), perform a self-loop walk for each encountered node and a single-edge quantum walk for each encountered edge.
@@
    Convert the sequence of the quantum walks to the circuit representation (which may include control reduction and CX forward- and backward-propagation, as described below).
Frontier-based graph traversal order selects the next node to visit from a set of nodes adjacent to already visited nodes. Many different heuristics can be used to construct the sequence of walks in step 1 or choose a particular traversal order in step 2. Initially, the system starts with all amplitude in the root state, which can be constructed from the ground |$\ket{0}^{\otimes n}$ state by applying the X gates to the necessary qubits.
An arbitrary amount of amplitude (absolute value) from the source can be transferred to the destination for any pair of the connected states corresponding to each single-edge walk, and all non-zero amplitude states are connected transitively via a tree. Therefore, an arbitrary distribution of absolute values of amplitude (but not phases) can be established on the set of non-zero amplitude basis states via single-edge walks.
Each self-loop walk on a basis state |$\ket{z_{i}}$ establishes the correct phase of the corresponding coefficient $c_{i}$, but does not change the absolute value of it. Therefore, performing self-loop walks on each non-zero amplitude basis state establishes the correct phases of the corresponding coefficients, without interfering with the absolute values established by the single-edge walk. Therefore, an arbitrary quantum state can be prepared with a sequence of single-edge walks and self-loop walks, as described in the algorithm above.
In the context of the state preparation, CRz gates can easily be used instead of CP gates to implement the self-loop walks, since every single-edge walk transfers the amplitude to a new basis state with zero amplitude. Thus, as described in the previous section, the same basis state can be used to establish the correct phase on the source state for the single-edge walk without additional CX conjugations. However, this will not apply to the leaf nodes (degree one vertices) in the tree, which will require another zero-amplitude state to interact with.
@@ IV.2 Control Reduction
Let $n$ denote the number of qubits and $S=\{\ket{z_{1}},\ket{z_{2}},...,\ket{z_{k}}\}$ denote the set of visited nodes. Suppose we wish to perform a single-edge walk from |$\ket{z_{j}}\in S$ to |$\ket{z_{\ell}}\notin S$ and |$\ket{z_{j}}$ and |$\ket{z_{\ell}}$ have a Hamming distance of 1. Let $b$ denote the qubit where |$\ket{z_{j}}$ is different from |$\ket{z_{\ell}}$, i.e. $z_{j}[b]\neq z_{\ell}[b].$ Let $D=\{\{k\,|\,z_{i}[k]\neq z_{j}[k],k\neq b\}\,|\,z_{i}\in S\setminus z_{j}\}$ denote the set of differing bits of the visited nodes with $z_{j}$ excluding the qubit $b$. For the gate-based representation of a single-edge walk from basis state |$\ket{z_{j}}$ to |$\ket{z_{\ell}}$, it is sufficient to control the CRz or CRx gate on any hitting set of $D$, where the values of controls are equal to the corresponding bits of $z_{j}$. A hitting set of $D$ is a collection of elements $h$ such that $h\cap d_{i}\neq\emptyset$ for all $d_{i}\in D$.
@@
When $z_{j}$ and $z_{\ell}$ have a Hamming distance greater than 1, we first have to update $S\rightarrow\tilde{S}$ with the conjugating CX gates, as described in Sec. III.1. Then we apply the previously described control reduction method to $\tilde{S}$. Note that in this case, any of the differing qubits can be used as a target for the CRz or CRx gates, with the appropriate CX conjugation. However, the target qubit cannot be chosen as a control, therefore the choice of target affects the efficiency of control reduction. In the following sections, when a target qubit is not explicitly dictated by a particular walk order, we consider all options for the target and choose the one that results in the minimum number of controls.
In order to minimize the number of controls, one needs to minimize the size of the hitting set, i.e. solve the minimum hitting set (MHS) problem, which is known to be NP-complete. As such, no polynomial algorithm is known to solve it exactly, but a number of heuristics exist that can provide good suboptimal solutions quickly [54, 55]. In practice, solving this problem for each walk drastically reduces the number of control qubits on many gates in the sequence (see Fig. 5). The MHS problems that we consider in the following sections are solved approximately.
Example: Consider the sequence of walks [[001, 111], [111, 110]]. Suppose we have implemented the first walk and are now implementing the second walk. The second walk transfers amplitude from 111 to 110. The current state is a superpostion of |$\ket{001}$ and |$\ket{111}$. Thus, we want the walk to affect only the amplitude of basis state 111 and leave 001 alone. Therefore, we need the controls for the Rx gate to be satisfied only by 111. Controlling the CRx gate on qubit 0 accomplishes this.
@@
IV.3 CX Forward- and Backward-Propagation
As described in Sec. III.1, single-edge quantum walks between states with Hamming distance greater than one require conjugation with CX gates on both sides of the CRx gate. The CX gates on the left side bring the desired basis states to a Hamming distance of 1, where they interact via the CRx gate, and the ones on the right move the basis states back to their original positions. Rather than immediately moving the states back after the application of the CRx gate, one can instead keep working with the transformed basis and move the right branch of the conjugating CX gates to the end of the circuit, which allows many of those CX gates from different single-edge walks to cancel. This is called forward-propagation of CX gates. The backward-propagation of CX gates is the same method applied to the reverse of the circuit. The advantage of the backward-propagation is that the CX gates are moved to the beginning of the circuit. Since the initial state is the ground state, these CX gates act trivially and can be removed.
Example: Consider the sequence of walks [[001, 111], [011, 110]]. The first walk [001, 111] has a Hamming distance of 2. For target qubit 1, we require CX10 to bring the states to a Hamming distance of 1. Applying this to the basis states we get [001, 011], [111, 010]. After applying the multicontrol gate(s) for the first walk, we do not rotate the basis states back to the original frame, but instead keep track of the CX rotation which will be applied at the end of the circuit. This is an example of CX forward-propagation.
IV.4 Leaf Node Considerations
When a walk is performed between two non-leaf nodes of the tree (see Fig. 4), we use a CRz gate to adjust the phase on the walk origin node, followed by a CRx gate to transfer the amplitude to the destination node, both of which have the same control set, since CRz operation does not change the number of non-zero amplitude basis states. The phase on the destination node is adjusted once we consider the first walk out of that node. However, when the destination node is a terminal (leaf) node, there are no other walks that start from it and we still need to fix its phase.
One way to solve this issue is to use a separate CRz gate, but it adds another multi-controlled gate to the circuit and the number of controls for it is likely to be greater compared to the walk into the leaf node due to the fact that we will need to distinguish the leaf from a larger number of non-zero amplitude states and control reduction will be less efficient.
An alternative approach is to combine this last CRz gate with the CRz and CRx gates from the walk leading into the leaf, as one gate that simultaneously transfers the probability amplitude and adjusts the phase on both ends of the walk. This reduces the number of multi-controlled operations from 3 to 1, but the downside of this approach is that the resulting gate will not belong to SU(2) group. As such, the number of CX gates in its decomposition will be proportional to $\mathcal{O}(n^{2})$, instead of $\mathcal{O}(n)$ necessary for the regular SU(2) gates, such as CRz or CRx. [56] Asymptotically, it makes this approach less efficient for sufficiently large values of $n$, but for our dataset it is still favorable.
IV.5 Tree Construction and Walk Order
Any tree constructed in the first step is, in principle, sufficient for state preparation, but not all trees result in the same CX gate count. Moreover, different walk orders on the same tree can can result in different number of CX gates as well. In this section we consider some of the options for tree construction and walk order.
Best Orders. We first describe heuristics that typically outperform state-of-the-art methods on average.
MHS Nonlinear. As demonstrated in section Sec. IV.2, control reduction is a powerful technique that allows to significantly simplify the state preparation circuit. The efficiency of it depends on how easily a particular state can be separated from the rest of the currently existing states with non-zero amplitudes, which in turn depends on the local density of states around the state that we want to separate. The basic idea of the following walk order heuristic is that to maximize the efficiency of control reduction, we need to walk through hard-to-separate states first, before the density of states around them increases, which allows us to separate them using smaller number of controls. Since the target qubit has a significant effect on the CX count, it is determined during the walk order construction as well.
More specifically, we construct a sequence of walks in reverse time order, from last to first. Given a list of basis states, we choose the state that requires the smallest number of controls to differentiate it from the rest of the basis states by solving the Minimum Hitting Set (MHS) problem. When two states are equally easy to distinguish, we choose the state with the largest number of bits that are different from the rest. Let us call this state $z_{1}$. The target qubit $t$ is the MHS bit with the smallest frequency. The set differentiated by $t$ is then ordered by the easiest element to distinguish via the MHS. When two elements are equally easy to distinguish, we order them by Hamming distance from $z_{1}$. The easiest element is set as $z_{2}$. The ($z_{1}$, $z_{2}$, $t$) that requires the fewest controls is added to the path. The number of controls required is upper bounded by the size of the MHS for $z_{1}$ plus the size of the MHS for $z_{2}$.
All the elements are then updated by the back-propagation of CX gates. $z_{2}$ is removed from the list and the process is repeated until every node is visited by the path. The circuit is constructed from the path of CTQWs in terms of the original basis states and target qubits and with CX backward-propagation (see Alg. 2 for pseudocode).
MHS Linear. This protocol follows the same procedure as MHS Nonlinear except that each step we assign $z_{1}$ of the previous pair as the $z_{2}$ for the next pair. This results in a linear graph. In contrast, in MHS Noninear $z_{2}$ of the next pair is chosen using the MHS methods.
Greedy Insertion. The linear greedy algorithm takes an ordered list of basis states and sequentially finds the best position to insert the next basis state. The classical complexity of this algorithm is $\mathcal{O}(nm^{4}+C)$, where $C$ is the classical complexity of the method that generates the initial ordering. Typically, the greedy insertion performs better when given an initial ordering that already performs well (see Alg. 3 for pseudocode).
Other Orders. The following are other orders we considered that on average typically do not outperform state-of-the-art methods.